# Heterogeneous Treatment Effects and

### v2

@@ 7 Conclusion
Social scientists routinely estimate HTEs with respect to a pre-treatment covariate with the stated intent of mechanism detection. We show that learning about the the activation of a mechanism using HTEs is far less straightforward than conveyed by current practice. Specifically, any link between a covariate (moderator) and a mechanism requires an exclusion assumption, so that covariate does not moderate other effects of treatment on an outcome. Even when this assumption holds, we can only use HTEs to affirm the activation of a mechanism when (1) HTEs exist and (2) the outcome is directly affected by a mechanism. Outside this case, HTEs do not provide sufficient information to show that a mechanism is active or inactive. In this sense, HTEs analysis should not be used to “rule out” mechanisms (or show that they are inert).
While mechanism detection is presently the modal use of HTEs in recent work in political science (see Table 1), it is not the only use of HTE. Our results speak to contexts where mechanistic analysis is the current aim. In current practice, HTEs are also increasingly used for extrapolation of treatment effects to different populations/settings (Egami and Hartman, 2022; Devaux and Egami, 2022) and the targeting of treatments (Athey, Tibshirani, and Wager, 2019; Kitagawa and Tetenov, 2018). Our work does not undermine these applications of HTEs, because these applications do not rely on questions of how to attribute observed effects to mechanisms (Slough and Tyson, 2024a).
Our analysis raises a number of issues and opportunities for future research to build upon. In particular, we emphasize the need to distinguish between the level at which mechanisms operate (e.g., on utility) and the outcomes we observe. This distinction has underappreciated implications for multiple quantitative methods to detect mechanism activation, including mediation analysis and investigation of the sign of treatment effects. Our framework can help to clarify the relationship between causal mechanisms and other applications of HTEs to clarify the (applied) theoretical foundations of these approaches.
@@ Appendix F Proofs of Propositions
We prove propositions under general multi-mechanism cases, as stated in the C. All propositions in the main text are all re-stated in the special case of one mechanism. Assumption 3 is only invoked in the multiple mechanism generalization.
@@ F.1 Proof of Proposition 1
Equation (36) follows from the linearity of expectations and the decomposition of the total effect in 31. Equation (37) is guaranteed by the linearity of $L$.
@@ F.3 Proof of Proposition 3
    (48) in the above theorem should be understood as
@@ F.4 Extension of Proposition 4
Here, will assume $y_{i}\in\mathbb{R}$ in (50) has no substantive interpretation. In practice, values of $y_{i}$ are typically normalizations, that are arbitrarily determined by the researcher. As such, the value is independent of model parameters.
@@ G.1 Discrete Outcomes under Monotonicity Assumptions
We ask whether researchers can differentiate these two cases when motonicity holds for the first DGP (e.g., an assumption of monotonicity). First, given (56) and (57), note that:
@@
If, under certain $x$ and $z$, (60) and the first term of (59) have the same sign and the second term of (59) has the opposite sign and is sufficiently large, (60) and (59) will have different signs. Moreover, If $e$ is uniformly distributed, then $f^{\prime}_{e}=0$ and thus equation (60) is equal to 0 while equation (59) is non-zero.
@@ G.2 Assumptions on the Latent Utility Distributions
$F(\varepsilon^{i})=\frac{\varepsilon^{i}}{1+\varepsilon^{i}}$

### v3

@@ 7 Conclusion
Social scientists routinely estimate HTEs with the goal of understanding which mechanisms generate treatment effects. By providing the first theoretical analysis of the relationship between HTEs and mechanisms, we show that detecting mechanisms with HTEs is far less straightforward than implied by current practice. Specifically, any link between a covariate (moderator) and a mechanism requires an exclusion assumption, so that covariate does not moderate other effects of treatment on an outcome. Even when this assumption holds, we can only use HTEs to affirm the activation of a mechanism when (1) HTEs exist and (2) the outcome is directly affected by a mechanism. Outside this case, HTEs do not provide sufficient information to show that a mechanism is active or inactive. In this sense, HTEs analysis should not be used to “rule out” activation of a mechanism.
At present, mechanism detection is the modal use of HTEs in political science (see Table 1) and the modal method for mechanism detection (Blackwell, Ma, and Opacic, 2024). However, mechanism detection is not the only use of HTE. Our results speak to contexts where mechanistic analysis is a goal. HTEs are also increasingly used for extrapolation of treatment effects to different populations/settings (Egami and Hartman, 2022; Devaux and Egami, 2022) and the targeting of treatments (Athey, Tibshirani, and Wager, 2019; Kitagawa and Tetenov, 2018). Our work does not directly speak to these uses of HTEs, because these methods do not seek to attribute observed effects to mechanisms (Slough and Tyson, 2024a).
Our analysis raises a number of issues and opportunities for future research to build upon. In particular, we emphasize the need to distinguish between the level at which mechanisms operate (e.g., on utility) and the outcomes we observe. This distinction has underappreciated implications for multiple quantitative methods to detect mechanism activation, including mediation analysis, analysis of treatment effects on intermediate outcomes, and efforts to link the sign of treatment effect to a (set of) mechanism(s). One potentially fruitful avenue for continued use of HTEs for mechanism detection would be to move from the presence of HTEs to their magnitude, as we outline in Appendix G.3. Adoption of this approach will rely on the the development of closer links between theoretical mechanisms and reduced-form treatment effects than is current practice.
@@ Appendix F Proofs of Propositions
We prove propositions under general multi-mechanism cases, as stated in the Appendix C. All propositions in the main text are all re-stated in the special case of one mechanism. Assumption 3 is only invoked in the multiple mechanism generalization.
@@ F.1 Proof of Proposition 1
Equation (36) follows from the linearity of expectations and the decomposition of the total effect in 11. Equation (37) is guaranteed by the linearity of $L$.
@@ F.3 Proof of Proposition 3
    Equation (48) in the above theorem should be understood as
@@ F.4 Extension of Proposition 4
Here, will assume $y_{i}\in\mathbb{R}$ in equation (50) has no substantive interpretation. In practice, values of $y_{i}$ are typically normalizations, that are arbitrarily determined by the researcher. As such, the value is independent of model parameters.
@@ G.1 Discrete Outcomes under Monotonicity Assumptions
We ask whether researchers can differentiate these two cases when motonicity holds for the first DGP (e.g., an assumption of monotonicity). First, given equation (56) and equation (57), note that:
@@
If, under certain $x$ and $z$, equation (60) and the first term of equation (59) have the same sign and the second term of equation(59) has the opposite sign and is sufficiently large, equation (60) and equation (59) will have different signs. Moreover, If $e$ is uniformly distributed, then $f^{\prime}_{e}=0$ and thus equation (60) is equal to 0 while equation (59) is non-zero.
@@ G.2 Assumptions on the Latent Utility Distributions
$F(\varepsilon^{i})=\frac{e^{\varepsilon^{i}}}{1+e^{\varepsilon^{i}}}$
@@
G.3 Assumptions about the Magnitude of HTEs
In practice, the exclusion assumptions may be too restrictive (strong). Consider the scenario depicted in the figure below, where our focus is on the activation of mechanism $M_{1}$. However, the covariate $X_{k}$ may influence mechanism $M_{2}$ and/or the direct effect. Our results indicate that we cannot learn about the activation of $M_{1}$ from the existence of HTE in $X_{k}$ since the exclusion assumption is violated. Here, we consider what might be learned from observing a large (in magnitude) HTE. This intuition can be formalized using Bayesian inference.
[…]
Figure A.7: The blue and yellow dashed arrows represent potential exclusion assumption violations.
Suppose that we have a prior $p$ over the activation of mechanism $M_{1}$. We characterize the violation of the exclusion assumptions—e.g., HTEs (the magnitude of the difference of CATEs) in $X_{k}$ that are generated by other mechanisms—with a normal distribution, $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$. The magnitude $|\mu_{2}|$ captures the severity of the violation. We assume that the covariate $X_{k}$ generates greater heterogeneity via the mechanism of interest, $M_{1}$, which we capture with the normal distribution $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$, where $|\mu_{1}|>|\mu_{2}|$. Then, the posterior after we observe such a large HTE is given by
$\mathbb{P}[M_{1}\text{active}\mid\text{as extreme as HTE}]$ $=\frac{p\mathbb{P}[\text{HTE}|\text{active}]}{p\mathbb{P}[\text{HTE}|\text{active}]+(1-p)\mathbb{P}[\text{HTE}|\text{not active}]}$
$=\frac{p\Phi(-2|\frac{HTE-\mu_{1}-\mu_{2}}{\sigma_{1}+\sigma_{2}}|)}{p\Phi(-2|\frac{HTE-\mu_{1}-\mu_{2}}{\sigma_{1}+\sigma_{2}}|)+(1-p)\Phi(-2|\frac{HTE-\mu_{2}}{\sigma_{2}}|)}$
where $\Phi(\cdot)$ is the CDF of standard normal distribution.
A Bayesian analyst might specify a prior distribution for the activation of a mechanism, denoted as $p\in[0,1]$, using a Beta distribution, $Beta(a,b)$. Subsequently, based on observed data, the posterior density of $p$ is
$\mathbb{P}[p|\text{HTE}]=\frac{dBeta(a,b)[\theta\Phi(-2|\frac{HTE-\mu_{1}-\mu_{2}}{\sigma_{1}+\sigma_{2}}|)+(1-\theta)\Phi(-2|\frac{HTE-\mu_{2}}{\sigma_{2}}|)]}{\int_{0}^{1}dBeta(a,b)[\theta\Phi(-2|\frac{HTE-\mu_{1}-\mu_{2}}{\sigma_{1}+\sigma_{2}}|)+(1-\theta)\Phi(-2|\frac{HTE-\mu_{2}}{\sigma_{2}}|)]dp.}$
For readers well-versed in Bayesian statistics, the posterior distribution of $p$ can be characterized using a variety of sampling techniques. For less familiar readers, we have provided a straightforward R script, which requires only the ggplot2 package and is designed to be user-friendly. It automatically generates figures like Figure A.8 to illustrate both the prior and posterior distributions, where we set $HTE=5$,$\mu_{1}=1$, $\mu_{2}=0.5$, $Beta(2,3)$, and $\sigma^{2}_{1}=\sigma^{2}_{2}=1$ below.
Figure A.8: Illustration of prior and posterior beliefs under a model in which $HTE=5$,$\mu_{1}=1$, $\mu_{2}=0.5$, $Beta(2,3)$, and $\sigma^{2}_{1}=\sigma^{2}_{2}=1$. The densities reflect the prior and posterior distributions of $p$, the probability that mechanism $1$ is active. The shift in the distribution to the right from prior to posterior suggests that the large HTE strengthens our belief that mechanism $1$ is active.
The strategy described above can also be applied to conduct sensitivity analysis. For instance, starting with a non-informative prior, one can determine how large $\mu_{2}$ needs to be—measured as a proportion of the total Average Treatment Effect (ATE)—to result in a posterior probability of the mechanism being active that is less than $50\%$.